OpenStudy (mathhelp346):

Answer and write in scientific notation: (9X10^5)-(3X10^5)

OpenStudy (anonymous):

\[3*10^5*(3-10^{-8})\]

OpenStudy (kropot72):

I see you edited your question. Th answer is as follows: \[(9\times10^{5})-(3\times10^{5})=10^{5}(9-3)=6\times10^{5}\]

OpenStudy (mathhelp346):

what about the other 10^5

OpenStudy (kropot72):

\[10^{5}\]is a common factor to the two bracketed terms in the question. Therefore it was taken outside a bracket leaving just a simple subtraction. Have you checked out the answer?

OpenStudy (mathhelp346):

? wouldnt it be 10^10 though

OpenStudy (kropot72):

Did you mean to multiply or subtract the two bracketed terms in the question?

OpenStudy (mathhelp346):

idk?

OpenStudy (mathhelp346):

but if u subtracted 10^5 and 10^5 would that be 0??

OpenStudy (kropot72):

You are right that 10^5-10^5=0. But that is not the original question which shows 3 times 10^5 subtracted from 9 times 10^5.

OpenStudy (mathhelp346):

so you need 9-3 which gave u 6 right

OpenStudy (mathhelp346):

and then u did 10^5-10^5?

OpenStudy (kropot72):

If the question is (9*10^5)-(3*10^5) then first take out the common factor 10^5 giving: 10^5(9-3) and do the subtraction to give10^5*6. Then express 10^5*6 in scientific notation as 6.0 * 10^5

OpenStudy (mathhelp346):

ok ty